# FRIDA to the Rescue! Analyzing Synthetic Data Effectiveness in Object-Based Common Sense Reasoning for Disaster Response

### v1

@@ Abstract
Large Language Models (LLMs) have the potential for substantial common sense reasoning. However, these capabilities are often emergent in larger models. This means smaller models that can be run locally are less helpful and capable with respect to certain reasoning tasks. To meet our problem space requirements, we fine-tune smaller LLMs to disaster domains, as these domains involve complex and low-frequency physical common sense knowledge. We introduce a pipeline to create Field Ready Instruction Decoding Agent (FRIDA) models, where domain experts and linguists combine their knowledge to make high-quality seed data that is used to generate synthetic data for fine-tuning. We create a set of 130 seed instructions for synthetic generation, a synthetic dataset of 25000 instructions, and 119 evaluation instructions relating to both general and earthquake-specific object affordances. We fine-tune several LLaMa and Mistral instruction-tuned models and find that FRIDA models outperform their base models at a variety of sizes. We then run an ablation study to understand which kinds of synthetic data most affect performance and find that training physical state and object function common sense knowledge alone improves over FRIDA models trained on all data. We conclude that the FRIDA pipeline is capable of instilling general common sense, but needs to be augmented with information retrieval for specific domain knowledge.
@@ 1 Introduction
Which of the following would be most dangerous if it collapsed? This question is fairly trivial for humans to answer, but requires several types of semantic knowledge. First, one must know what objects are capable of collapsing. One must also know the general size of these items and the item’s other functions to assess the danger. A collapse is also a change of state that fundamentally shifts the use of these objects; a collapsed chair is more likely to cut, scrape, or be carried if the chair folds. All of this knowledge is needed to answer this question, and all of it is embedded in our semantic understanding of things that can cause danger and things that can collapse.
@@
As LLMs have improved exponentially, their abilities at reasoning about objects have improved as well. LLMs have long proven to encode physical world knowledge Petroni et al. (2019), and their embeddings can improve physical understanding of an environment and its objects both within and beyond a fine-tuned domain Cohen et al. (2024). However, much of this improvement only emerges in larger models trained on more data Wei et al. (2022a). This makes these essential semantic capabilities less accessible, particularly for running LLMs locally on edge devices. We thus wanted to answer: how can we imbue all the physical common sense and semantics needed for smaller models to be more capable at understanding the physical world?
To answer this question, we turned to improving small LLMs in the disaster relief domain. The first reason for this was because disasters require a lot of both general and domain-specific semantic knowledge: one needs to know both the standard objects in daily life and the specialized tools required for search and rescue, such as hydraulic pumps and concrete saws. Secondly, the specific knowledge (and to a lesser extent, the general knowledge) needed for reasoning varies by disaster, which requires a high level of adaptability in any potential pipeline. Finally, most publicly available data on disasters is social media-based reactions Godinho (2024), which does not pertain much to our subdomain of the objects used during these events. This is due to the writer’s assumption of latent semantic knowledge (e.g., everyone knows that counters don’t collapse, so no one makes a point to tweet that) Gordon and Van Durme (2013). This makes the task more novel for LLMs which likely have not seen explicit reasoning in this domain during pre-training.
@@
We present a pipeline to create Field Research and Instruction Decoding Agent (FRIDA) models. For FRIDA, we leveraged both disaster and linguistic expertise to create gold-standard instructions that in turn are used as a basis for synthetic data generation, as seen in Figure 2. This synthetic data is then used to fine-tune smaller models to increase Like its rescue dog eponym, [^1] https://en.wikipedia.org/wiki/Frida_(dog) our FRIDA models were initially developed and tested for earthquake disaster relief, based on expert knowledge pertaining to the February 6th, 2023 earthquakes in Turkiye and Syria Arranz et al. (2023).
@@
We found our FRIDA models out-performed their base models overall in both exact match and embedding vector similarity Aynetdinov and Akbik (2024), regardless of model size or architecture. Knowing that our pipeline improved performance, we wanted to investigate which synthetic data were most influential in that improvement. Do accomplish this, we ran an ablation study where we fine-tuned a variety of small LLMs on subsets of our synthetic data corresponding to a specific type of object-based common sense. We call these resulting models the ablation FRIDA (aFRIDA) models. We found that aFRIDA models trained on general common sense significantly outperformed models trained on the domain-specific common sense knowledge and the full FRIDA models themselves. We posit that FRIDA succeeds in improving object-related general common sense and is a strong basis for improving domain-specific general common sense as well.
@@
    An expert-in-the-loop pipeline (Figure 2) for generating specific and high-quality synthetic data that can be used for fine-tuning when man-made data are not feasible to obtain, as well as the resulting gold-standard datasets
@@
    A synthetic dataset of 25,000 instructions relating to common sense and earthquake with accompanying analysis
@@
    The FRIDA 1B, 3B, Minstal 8B, and LLaMa 8B models, trained on the above synthetic data.
@@
    A series of ablation FRIDA (aFRIDA) models trained on subsets of the synthetic dataset
An anonymous github containing code and a complete example of the FRIDA pipeline is currently available. [^2] https://anonymous.4open.science/r/FRIDA-0215-D02A/
Figure 2: The pipeline to create the FRIDA suite of models. A search and rescue expert fills out a survey on the relevant tasks and objects a robot needs to be aware of, then a semantics expert adds those terms to the ontology and fills in the templates to generate new seed instructions for a variety of different disasters. These seed sentences are then utilized to generate synthetic data for fine-tuning an LLM with the necessary expertise on the specific disaster at hand.
2 Background
@@ 2.1 Synthetic Data Generation
This paper is inspired in particular by the pipeline used in Wang et al. (2022). Wang et al., 2022 hand crafted 175 instructions that were used for 8-shot learning to prompt GPT’s text-davinci-001 model to generate more than 50,000 instructions for a generic and ungrounded AI assistant. These instructions were then used to fine-tune text-davinci-001. The authors found that their method and resulting fine-tuned model performed comparably to OpenAI’s GPTInstruct Wang et al. (2022). Taori et al., 2023 innovated on Wang et al., 2022 by fine-tuning a separate, smaller language model with a different architecture, as opposed to fine-tuning the same model that generated the data. They subsequently found that their resulting model’s answers were rated as highly as GPT’s davinci-text-003. We adopt the latter innovation in our approach, wherein we leverage Gemini 1.5 Flash for generating synthetic data, which we use to fine-tune LLaMa3 and Minstal instruction-tuned models.
2.2 Data Creation
We developed an expert-in-the loop pipeline to generate high-quality seed data that leverages disaster-relief and linguistic expertise. The purpose of this pipeline is to enable quick and efficient fine-tuning of small LLMs capable of addressing critical informational needs in specific disasters. The details of the pipeline are described in a previous work, here we provide a brief overview. We developed a series of templates that can be filled in with vocabulary from an affordance ontology, based on the Rich Event Ontology Kazeminejad et al. (2018). This affordance ontology is extended to serve as an ontology of disaster-related objects and their functionalities, as defined by the objects’ PropBank semantic roles labels Palmer et al. (2005).
To fill in these templates with proper data, a disaster expert first provides information about the relevant objects and situations a they would encounter in their work. For this paper, the authors simulated this step by gathering existing resources authored by experts on the Turkiye-Syria Earthquake recovery efforts Arranz et al. (2023). After gathering domain-specific data, linguists go through a template-filling pipeline. Summarily, the linguists select the relevant vocabulary from the expert knowledge to add to the aforementioned affordance ontology. They then use this ontology and template-specific generation instructions to fill in these templates to create gold-standard data, our “seed instructions”. These templates are formatted as multiple choice questions with semantically distinct answers. Some examples of this process, as well as some synthetic instructions that result from 5-shot prompting, can be seen in Table 1.
@@ 2.2 Data Creation
Gold Standard “Seed” Instruction What state should a drawbridge be in for cars to cross a river? A) Lowered or B) Raised
@@
Gold Standard “Seed” Instruction What role do hydraulic lifts play in rescuing people after an earthquake?
@@
Table 1: 2 Examples of templates and their corresponding gold standard and synthetic instructions. Note that the blanks in the first template can only be filled in by objects with multiple states (linguistic knowledge), while the blanks in the second template can only be filled in with specific tools (disaster expert knowledge).
In total, we created 26 templates, which were then grouped by topic or type of common sense. For example, the first example in Table 1 describes a change in state, which is often a change in the shape, size, or weight of an object. It therefore was put in the category Relative Sizes, the category concerned with how the object exists in space. The second example in table 1 describes how a specific tool related to the disaster is used. It was placed in the category Required Equipment, the category concerned with how to use the specific tools required for the disaster. A complete list of categories, the templates within them, and examples, can be found in the Appendix Table 4.
3 Methods
@@ 3.1 Synthetic Dataset Generation and Analysis
Metric Range Averages
Instruction length 19.11-36.98 39.44 words
Maximum ROUGE 0.51-0.73 0.67
Reasonableness 0.80
Informativeness 0.49
@@
Table 2: The top table contains overall statistics for our synthetic dataset. There is a large range in values across categories when compared to the average for both length and ROUGE values. We had high average human ratings of reasonableness to fitting the prompt, but only fair informativeness. The bottom table is the sizes of the training and development datasets used for fine-tuning FRIDA and its ablations.
The dataset we use in this work is about the Turkiye-Syria Earthquake. Our data is formatted as a user-assistant chat in order to more closely align our data to the general instruction tuning all our base models received. In addition to the seed data, we also developed a separate gold-standard evaluation dataset. The final result is that each template has 5 seed instructions for synthetic data generation (130 total instructions) and at least 4 evaluation instructions (119 examples).
For each template, we used 5-shot prompting with Gemini-1.5-flash to generate 980 examples of said template Team (2024a). We chose Gemini for its accessible and affordable API, as well as its high scores on our evaluation (0.725 exact match accuracy, 0.94 average Semscore, see section 3.3). We prompted Gemini to return 40 instructions per API call. To ensure our synthetic data were robust, we used ROUGE scoring Lin (2004) to ensure Gemini was not giving us duplicates of previously generated instructions. Depending on the template, the cut-off ROUGE score went from 0.8 for templates with more varied language to 0.97 for templates with very structured wording. We also increased model temperature for the more structured templates to increase diversity of responses.
We get a sense of the resulting synthetic dataset from the topmost chart in table 2. We automatically evaluated for instruction length and each instruction’s maximum ROUGE score with the other instructions in the dataset. We found we had substantial average instruction length, and reasonable ROUGE scores given our data is template-based. However, there was a large range in both metrics across the different template categories. We attribute this to the overall complexity of the individual templates, which vary substantially.
We then randomly selected 190 synthetic instructions, with 4-5 examples from each template and had 2 authors examine them. The authors rated these instructions on 2 binary metrics: reasonableness and informativeness. An instruction is marked as “Reasonable” if the model sufficiently followed the template’s general format and intent. An instruction marked as “Informative” means the question is non-trivial. For instance, the instruction “Which would be most likely to cause a serious injury if it fell from a roof? A leaf, a small pebble, a large tile, a piece of paper, or a feather” follows the injury template “Which of the following would be most likely to cause an injury in [SCENARIO]?”. However, the four wrong answers are so harmless the instruction becomes too easy to be meaningful. It follows that an instruction can only be informative if it is reasonable. As seen in table 2, 80% of sampled instructions were reasonable, but only 50% were informative. The 2 authors had Cohen’s $\kappa$ values of 0.28 for reasonableness and 0.35 for informativeness, meaning the authors had fair agreement for both metrics.
@@ 3.2 FRIDA Model construction
We used our synthetic dataset to fine-tune the 1 Billion, 3 Billion, and 8 Billion parameter Instruct models from the LLaMa-3 herd Team (2024b) as well as the Ministral 8B Instruction tuned model Team (2024c). We chose to use the LLaMa suite due to its strong performance on its open source small LLMs, as well as it having multiple small instruction tuned models of different sizes Team (2024b). We chose Ministral 8B to serve as a comparison, since it is trained with sliding window attention, unlike the LLaMa models Team (2024c). It also was released after the LLaMa 3 herd and outperformed it in many metrics Team (2024c). We chose to fine-tune the instruct variations of these models because our task is based in answering questions. Fine-tuning specifics can be found in Appendix section C.

### v2

@@ Abstract
During Human Robot Interactions in disaster relief scenarios, Large Language Models (LLMs) have the potential for substantial physical reasoning to assist in mission objectives. However, these capabilities are often found only in larger models, which are frequently not reasonable to deploy on robotic systems. To meet our problem space requirements, we introduce a dataset and pipeline to create Field Reasoning and Instruction Decoding Agent (FRIDA) models. In our pipeline, domain experts and linguists combine their knowledge to make high-quality few-shot prompts used to generate synthetic data for fine-tuning. We hand-curate datasets for this few-shot prompting and for evaluation to improve LLM reasoning on both general and disaster-specific objects. We concurrently run an ablation study to understand which kinds of synthetic data most affect performance. We fine-tune several small instruction-tuned models and find that ablated FRIDA models only trained on objects’ physical state and function data outperformed both the FRIDA models trained on all synthetic data and the base models in our customized evaluation. We demonstrate that the FRIDA pipeline is capable of instilling physical common sense with minimal data.
@@ 1 Introduction
Which of the following would be most dangerous if it collapsed? This question, as seen in Figure 1, is fairly trivial for humans to answer, but requires several types of semantic knowledge. First, one must know which objects are capable of collapsing. One must also know the general size of these items and the their other functions to fully assess the danger the item poses. A collapse is also a change of state that fundamentally shifts the use of these objects; a collapsed chair could be more likely to cut or scrape someone, but it could also mean the chair can now be carried if the chair folds. All of this knowledge is needed to answer this question, and all of it is embedded in our semantic understanding of objects that can cause danger and objects that can collapse, both intentionally and unintentionally.
The ability to reason about objects is especially important in the context of human-robot interaction in disaster relief scenarios Bonial et al. (2024). For example, during search and rescue after an earthquake, a robot needs to know how to navigate partially collapsed buildings and how to use the many tools required to free people from the rubble. However, using robots to aid in disaster relief introduces many constraints. Because of the destruction a disaster can wreak, consistent internet connectivity cannot be assumed. For human safety, robots must be handled via radio in a secure location. This low-bandwidth communication means limited image data can be transmitted to the handlers, which rules out remote piloting Bonial et al. (2024). We therefore need an autonomous system that can reason about its environment and the relief tasks required.
@@
As LLMs have improved dramatically, their abilities at semantic reasoning about objects have improved as well. LLMs have long been proven able to encode physical world knowledge Petroni et al. (2019), and their embeddings can improve physical understanding of an environment and its objects both within and beyond a fine-tuned domain Cohen et al. (2024).
However, much of this improvement is found only in larger models trained on more data Wei et al. (2022a); Kaplan et al. (2020). This makes these essential semantic capabilities less accessible to our use case. Our robot cannot rely on an internet connection to make API calls. We instead must utilize the robot’s limited on-board computing power, which can be as little as 16 GB of virtual RAM on an array of GPUs Osteen (2025). That amount of GPU RAM can only reasonably run inference on a 13 Billion parameter model given the heuristics described in Anthony et al. (2023). Furthermore, this heuristic assumes that our robot is not running other processes in parallel, which is fairly unreasonable. We thus wanted to answer: Given our constraints, how can we imbue all the physical common sense and semantics needed for smaller LLMs to be more capable at understanding a disaster environment?
To answer our research question, we first test the effectiveness of fine-tuning smaller models on disaster relief data. However, available data proved to be an additional constraint. Most publicly available data on disasters is social media-based reactions Godinho (2024), which do not pertain much to our subdomain of disaster relief efforts. Furthermore, the specific knowledge (and to a lesser extent, the general knowledge) required for each mission varies by disaster. For example, after an earthquake, a robot needs to find survivors, while after a chemical spill, a robot needs to sample the environment for hazardous materials. Therefore, we need a method for generating training data for specific disasters, and we need to evaluate which data are most effective at improving robot performance.
@@
We present a pipeline to create Field Reasoning and Instruction Decoding Agent (FRIDA) models as a proof of concept for LLM viability in the disaster relief domain. For FRIDA, we leveraged both disaster and linguistic expertise to create gold-standard instructions that, in turn, are used as a basis for synthetic data generation, as seen in Figure 2. These synthetic data are then used to fine-tune smaller models that fit our memory constraints. Like its rescue dog eponym, [^1] https://en.wikipedia.org/wiki/Frida_(dog) our FRIDA models were initially developed and tested for earthquake disaster relief, based on expert knowledge pertaining to the February 6th, 2023 earthquakes in Turkey and Syria Arranz et al. (2023). Thus, the resulting models are small enough to effectively operate onboard a robot and are fine-tuned on specialized and inexpensive data, satisfying all of our use case constraints.
@@
To investigate which synthetic data most influenced model performance, we ran an ablation study where we fine-tuned the same small LLMs on subsets of our synthetic data corresponding to specific types of object-based reasoning. We call these resulting models the ablated FRIDA (aFRIDA) models. We found that aFRIDA models trained on general semantics and physical common sense had stronger overall performances than models trained on only domain-specific knowledge. Additionally, the best performing aFRIDA models scored better than their corresponding base models and FRIDA models trained on the entire synthetic dataset. We posit that FRIDA succeeds in improving object-related general common sense, but that small LLMs struggle with disaster-specific equipment usage
@@
    An expert-in-the-loop pipeline (Figure 2) for generating specific and high-quality synthetic data that can be used for fine-tuning when man-made data are not feasible to obtain, as well as the resulting gold-standard datasets.
@@
    A synthetic dataset of 25,000 instructions relating to object reasoning and earthquake response with accompanying analysis.
@@
    The FRIDA model, fine-tuned on Mistral AI’s Ministral 8B model with the above synthetic data, which investigates small LLM potential.
@@
    A series of ablated FRIDA (aFRIDA) models trained on subsets of the synthetic dataset to investigate which synthetic data were most effective.
  5. 5.
    An in-depth analysis investigating the challenges of imbuing physical common sense and complex object reasoning into LLMs.
An anonymous github containing code and a complete walkthrough of the FRIDA pipeline are currently available. [^2] https://anonymous.4open.science/r/FRIDA-0215-D02A/
Figure 2: The pipeline to create the FRIDA suite of models. A search and rescue expert fills out a survey on the relevant tasks and objects used in disaster response, then a semantics expert adds those terms to the ontology and fills in the templates to generate new seed instructions for a variety of different disasters. These seed sentences are utilized to generate synthetic data for fine-tuning an LLM with the necessary expertise on the specific disaster.
2 Related Work
2.1 LLMs Reasoning about the World
There are a wide variety of methods for leveraging LLMs for reasoning in a physical environment based on Chain of Thought prompting Wei et al. (2022b). These include variants like re-prompting Raman et al. (2022), which prompts the LLM to regenerate a plan if certain criteria aren’t met at certain steps, or Tree of Thought Yao et al. (2023), which generates a tree of potential steps and evaluates each potential path via either a breadth-first or depth-first search.
There are also methods that allow the LLM to take in environmental feedback in response to its output. For Inner-Monologue Huang et al. (2023), the LLM is given the option to ask for more scene descriptors from a human handler, which it then incorporates into its prompts, improving task completion and decreasing hallucination. Another example is SayPlan Rana et al. (2023), which uses 3D scene plans to iterate on proposed strategies until an effective path is discovered. Xie and Zou (2024) get feedback from LLMs themselves by using a wide variety of LLM agents to do various sub-tasks for planning, including generating a general outline, using external tools to gain information, and evaluating which plan is best.
One resource for improving LLM understanding of an object’s functions, also known as the object’s affordances, is Adak et al. (2024), who curate a dataset of naturally occurring sentences and corresponding images. They then transform them into inference, probing, and masking tasks for LLMs and Visual Language Models (VLMs). Their evaluation shows that VLMs do not have straightforward understandings of object affordances, but few-shot fine-tuning improves LLM and VLM performance on identifying object affordances.
2.2 Disaster Work and Natural Language Processing
Godinho (2024) completed a systematic search and analysis of over 100 peer-reviewed papers relating to Natural Language Processing (NLP) tools being applied to disasters. 85 of the 107 papers found were for analyzing social media, and the majority of papers focused on sentiment analysis, text classification, and information extraction tasks. Both the data sources and NLP tasks do not have a clear parallel with our objective.
While robots have been successfully deployed in disaster relief missions, the current state of the art is a human tele-handler in complete control of the robot Chiou et al. (2022); Kanazawa et al. (2023). This puts all of the cognitive burden on said tele-handler, and does not allow for the re-tasking and pivoting required in such a high-stakes, fast changing scenario Bonial et al. (2024). To move the state of the art from tele-handling to human-robot dialogue, Lukin et al. (2024) provide a corpus of simulated dialogues in a disaster scenario that are annotated for semantic meaning, dialogue structure, and visual common ground. However, this corpus works with a robot with limited abilities and does not touch on creating a system to reason about a wide variety of objects and disasters.
@@ 2.3 Synthetic Data Generation
This paper is inspired in particular by the pipeline used in Wang et al. (2022), who hand crafted 175 “seed” instructions. These seed instructions were used for 8-shot prompting of GPT’s text-davinci-001 model to generate more than 50,000 instructions for a generic and ungrounded AI assistant. These synthetic instructions were then used to fine-tune text-davinci-001. The authors found that their method and resulting fine-tuned model performed comparably to OpenAI’s GPTInstruct Wang et al. (2022). Taori et al. (2023) innovated on Wang et al. (2022) by fine-tuning a separate, smaller language model with a different architecture, as opposed to fine-tuning on the same model that generated the data. They subsequently found that their resulting model’s answers were rated as highly as GPT’s text-davinci-003.
3 Methods
3.1 FRIDA Seed Data
We developed an expert-in-the-loop pipeline to generate high-quality seed data that leverage expertise on both disaster-relief and semantics. The purpose of this pipeline is to enable quick and efficient fine-tuning of small LLMs to be capable of critical reasoning in specific disaster environments. The details of this pipeline are described in a previous work, here we provide a brief overview. We developed a series of templates that can be filled in with vocabulary from an affordance ontology based on the Rich Event Ontology Kazeminejad et al. (2018). This affordance ontology is extended to serve as an ontology of disaster-related objects and their functionalities, as defined by the objects’ PropBank semantic roles labels Palmer et al. (2005).
To fill in these templates with proper data, a disaster expert first provides information about the relevant objects and situations encountered in their work. For this paper, the authors simulated this step by gathering existing resources authored by experts on the Turkey-Syria Earthquake recovery efforts Arranz et al. (2023). After gathering domain-specific data, linguists go through a template-filling pipeline. Summarily, the linguists select the relevant vocabulary from the expert knowledge to add to the aforementioned affordance ontology. They then use this ontology and template-specific generation instructions to fill in the templates to create “seed” instructions. These templates are formatted as multiple choice questions with semantically distinct answers. Some examples of this process, as well as some of the synthetic instructions that result, can be seen in Table 1.
@@ 3.1 FRIDA Seed Data
Seed Instruction What state should a drawbridge be in for cars to cross a river? A) Lowered or B) Raised
@@
Seed Instruction What role do hydraulic lifts play in rescuing people after an earthquake?
@@
Table 1: Two Examples of templates and their corresponding gold standard and synthetic instructions. Note that the blanks in the first template can only be filled in by objects with multiple states (i.e. linguistic knowledge), while the blanks in the second template can only be filled in with specific tools (i.e. disaster expert knowledge).
Although some related work leverages the same seed sentences used for generating synthetic data to also evaluate the data Wang et al. (2022), we developed a separate evaluation to ensure that our evaluation was not unfairly biased. The seed and evaluation instructions include multiple choice answers, enabling more efficient evaluation and comparison of models.
@@ 3.2 Synthetic Dataset Generation and Analysis
Figure 3: The distribution of the synthetic data’s instruction length (top) and maximum ROUGE score (bottom). Averages are shown as black dashed lines. Our high average instruction length and general distribution shows synthetic instructions are sufficiently complex, and our average over each instruction’s top ROUGE score shows the instructions are sufficiently unique for this to be a challenging task.
The dataset we use in this work focused on search and rescue operations in the aftermath of the Turkey-Syria Earthquake Arranz et al. (2023). We had 26 templates grouped into 8 categories based on the type of knowledge they query (for all categories and examples, see Table 4). For each template, we made 5 seed instructions for synthetic data generation (130 total instructions) and at least 4 evaluation instructions (119 examples).
For each template, we used its corresponding seed instructions for 5-shot prompting with Gemini-1.5-flash to generate 980 synthetic instructions based on the given template Team (2024a). We chose Gemini as our synthetic data generator for its accessible and affordable API, as well as its high scores on our evaluation (93.9% average Semscore, see section 3.4). We prompted Gemini to return 40 instructions per API call. To ensure our synthetic data were unique, we used ROUGE scoring Lin (2004) to ensure Gemini was not giving us duplicates of previously generated instructions. Depending on the template, the cut-off ROUGE score went from 0.8 for templates with more varied language to 0.97 for templates with very structured wording. We also increased model temperature for the more structured templates to increase diversity of responses.
We get a sense of the resulting synthetic dataset from the histograms in Figure 3. We automatically evaluated for instruction length and each instruction’s maximum pairwise ROUGE score. We found we had substantial average instruction length, and reasonable ROUGE scores given that our data are template-based. There was a large range in both metrics across the different template categories, which we attribute to the overall complexity of the individual templates. Some templates require short instructions with binary answers, while others have longer instructions where all answers are sentences.
@@
Table 2: The number of instructions in the training and development datasets used for fine-tuning FRIDA and its ablations.
@@ 3.3 FRIDA Model construction
We used our synthetic dataset to fine-tune the 1 Billion, 3 Billion, and 8 Billion parameter Instruct models from the LLaMa 3 herd Team (2024b) as well as the Mistral AI’s Ministral 8B Instruction tuned model Team (2024c). We chose to use the LLaMa suite due to it having multiple small instruction tuned models of different sizes, with decently strong performanceTeam (2024b). We chose Ministral 8B to serve as a comparison, since it is trained with sliding window attention, unlike the LLaMa models trained with full attentionTeam (2024c). Additionally, Ministral 8B was released after the LLaMa 3 herd and outperformed the LLaMa models on a variety of metrics Team (2024c). We chose to fine-tune the instruct variations of these models because our task is based in answering questions.
Of the four fine-tuned models, the strongest fine-tuned model performance on our evaluation was from models trained on Ministral 8B. We hypothesize that this is due to the architectural differences between Mistral AI and Meta AI models. Specifically, sliding attention could be helpful in focusing the model’s attention on the instruction content instead of the multiple choice answers. Additionally, Ministral 8B’s sliding attention mechanism is more memory and time efficient, making it more practical for deployment on a robot Team (2024c). As such, we focused our analysis on FRIDA and aFRIDA models based on Ministral 8B, since they are the most conceivable models to work in a robotic system in the near term. Results for the LLaMa models can be found in our anonymous github. Fine-tuning specifics can be found in Appendix section B.